OpenStudy (anonymous):

evaluate x if log2 (1+x) + log2 (5-x) - log2 (x-2)=3

OpenStudy (anonymous):

\[ \log_2 (1+x) + \log_2 (5-x) - \log_2 (x-2)=3\]\[ \log_2[~~ (-x^2+4x+5) \div (x-2)~~]=3\]\[ \log_2[~~ (-x^2+4x+5) \div (x-2)~~]=3~\log_22\]\[ \log_2[~~ (-x^2+4x+5) \div (x-2)~~]=\log_28\]\[ (-x^2+4x+5) \div (x-2)=8\]\[ -x^2+4x+5=8x-16\]\[ -x^2-4x+21=0\]

OpenStudy (anonymous):

\[ -x^2-4x+21=0\]\[x^2+4x-21=0\]\[(x+7)(x-3)=0\]

OpenStudy (anonymous):

thanks you very much , but now i have another question, can you help me? log5 (x)= 4logx (5) , evaluate x. is it i have to change both of the log into same base?

OpenStudy (mathmate):

@kY_Tan To complete the solution, you wukk need to validate each of the solutions. Whenever you're dealing with log functions, you need to verify that the solution does not give a negative argument for the log.

OpenStudy (mathmate):

*will

OpenStudy (anonymous):

@mathmate thanks

OpenStudy (mathmate):

@kY_Tan you're welcome! :)